# HardcoreLogic: Challenging Large Reasoning Models with Long-tail Logic Puzzle Games

### v2

@@ Abstract
Large Reasoning Models (LRMs) have demonstrated impressive performance on complex tasks, including logical puzzle games that require deriving solutions satisfying all constraints. However, whether they can flexibly apply appropriate rules to varying conditions, particularly when faced with non-canonical game variants, remains an open question. Existing corpora focus on popular puzzles like 9x9 Sudoku, risking overfitting to canonical formats and memorization of solution patterns, which can mask deficiencies in understanding novel rules or adapting strategies to new variants. To address this, we introduce HardcoreLogic, a challenging benchmark of over 5,000 puzzles across 10 games, designed to test the robustness of LRMs on the “long-tail” of logical games. HardcoreLogic systematically transforms canonical puzzles through three dimensions: Increased Complexity (IC), Uncommon Elements (UE), and Unsolvable Puzzles (UP), reducing reliance on shortcut memorization. Evaluations on a diverse set of LRMs reveal significant performance drops, even for models achieving top scores on existing benchmarks, indicating heavy reliance on memorized stereotypes. While increased complexity is the dominant source of difficulty, models also struggle with subtle rule variations that do not necessarily increase puzzle difficulty. Our systematic error analysis on solvable and unsolvable puzzles further highlights gaps in genuine reasoning. Overall, HardcoreLogic exposes the limitations of current LRMs and establishes a benchmark for advancing high-level logical reasoning.
Leaderboard: https://huggingface.co/spaces/JunsWan/HardcoreLogic
Dataset: https://huggingface.co/datasets/xhWu-fd/HardcoreLogic
Code: https://github.com/ljcleo/hardcore-logic
@@ 1 Introduction
Despite recent successes on benchmarks such as Enigmata (Chen et al., 2025a) and ZebraLogic (Lin et al., 2025a), whether LRMs are genuinely capable of true logical reasoning, i.e., flexibly apply appropriate rules to relevant conditions to derive correct conclusions, remains an important question. Take Sudoku as an example: while most real-world puzzles follow the canonical 9x9 format with nine 3x3 zones, variants with alternative constraints or irregular subgrids often prove challenging even for humans. Similarly, existing corpora exhibit a severe imbalance between canonical and non-canonical logic puzzles, making models prone to overfitting to canonical puzzles (Cohen-Inger et al., 2025), leading to difficulties in solving non-canonical variants that fall into the long-tail of the distribution. This limitation manifests in two specific ways: (1) Models recognize only the canonical form of logical puzzles; when given a variant, they either struggle in understanding the new rules or ignore them, leading to faulty reasoning. (2) Models develop fixed solution strategies and reasoning patterns to solve canonical puzzles; even when they successfully understand the variant, they still apply a mismatched solution strategy, eventually producing errors or being confused.
@@
We further conduct a systematic error analysis to probe the underlying causes of model failures on HardcoreLogic. For solvable puzzles, we classify erroneous responses into six categories, and find that factual errors dominate across models, while more powerful models tend to exhibit brute-force errors, attempting exhaustive searches rather than strategic reasoning. Besides, models’ misunderstanding of problem constraints and misapplication of rigid rules lead to significant performance drops. For unsolvable puzzles, our analysis reveals that models performing well on solvable problems genuinely recognize unsolvability better. However, weaker models like Minimax-M1 may output ”unsolvable” simply when they fail to find an answer, rather than through true recognition of logical unsatisfiability. When models fail to recognize unsolvability, we observe that stronger models mainly fail due to erroneous reasoning or inability to output answers within token budgets, while weaker models tend to force out solutions even without successfully deriving them. These highlight the need to improve models’ deep reasoning capabilities and robustness against degenerate behaviors. Overall, our contributions are threefold:
@@ 2.1 Preliminary: logical puzzle games
Logic game category
In HardcoreLogic, we focus on 10 types of logic games spanning 6 puzzle categories, including 8 challenging subtasks sourced from Enigmata (Chen et al., 2025a), the Zebralogic game from the ZebraLogic dataset (Chen et al., 2025a), and a classic Hanoi game synthesized by ourselves following its standard rules. All these three sources constitute the Original data used for comparison with HardcoreLogic. Specifically, HardcoreLogic covers the following 6 categories: (1) logic puzzle, (2) grid puzzle, (3) search puzzle, (4) pattern puzzle, (5) graph puzzleand (6) sequential puzzle. The 10 specific games are Zebralogic, Sudoku, Skyscraper, Kakurasu, Crypto, Navigation, Binario, Minesweeper, Hanoi and Hitori. A detailed introduction to each puzzle category is provide in Appendix B.1.
@@ 2.2 Long-tail transformation
Standard logic puzzles are constrained in size, form diversity, and rule design, and thus fail to capture the irregularity and scale of real-world reasoning. To systematically construct more challenging evaluation data, we introduce a set of long-tail transformations that extend puzzles along three orthogonal dimensions: Increased Complexity, Uncommon Element, and Unsolvable Puzzle.
@@ 2.3 Complexity analysis
To systematically evaluate the hardness introduced by our long-tail transformations, we conceptualize complexity as a four-dimensional construct: Search Space Expansion (IC1), Constraint Strengthening (IC2), Form Mutation (UE1), and Rule Mutation (UE2). Each transformation is associated with dedicated quantitative metrics, and we compare all generated puzzles against the original benchmark. In the following, we present quantitative analyses of these four transformation types to demonstrate how each contributes to increased puzzle difficulty.
@@
This dimension captures the growth of candidate assignments induced by empty cells. Closed-form formulas are derived for each puzzle family (Appendix). For instance, in Binario, if $N$ cells are empty , the search space is $\lvert S\rvert=2^{N}$. Figure 2 shows the average log-scale search space across five puzzle families, confirming that HardcoreLogic systematically enlarges the combinatorial space.
@@
  • For CSP-based puzzles (e.g., Zebralogic, Binario), we encode instances into Z3 and collect: (i) Decisions: Explicit branching steps made by the solver; (ii) Conflicts: Backtracking events where partial assignments lead to contradictions. Larger counts indicate more complex search spaces and stronger constraint interactions, reflecting higher difficulty.
@@
Form mutation (Ue1)
Form mutation introduces novel symbols or forms that preserve puzzle validity but complicate comprehension. Since symbolic solvers cannot capture this representational difficulty, we measure it using perplexity, the inverse probability assigned by a pretrained LRM, which quantifies how surprising an instance appears. Higher perplexity values indicate that mutated forms impose greater representational complexity for LRMs. Figure 3 (light orange background) presents boxplots comparing the perplexity distributions of Original and HardcoreLogic instances; form mutation consistently results in higher perplexity and thus greater representational difficulty.
Rule mutation (Ue2)
Rule mutation modifies or extends the logical rules governing puzzles, forcing solvers to adapt to new structural constraints.
@@
As shown in Figure 3 (light purple background), mutated-rule puzzles consistently yield higher solver statistics, indicating rule changes intensify reasoning complexity. A notable exception is the minesweeper dataset with ”landmine clusters”: numerical clues now represent adjacent landmine clusters, and more clues are added to ensure unique solutions—this reduces the search space, making required decisions lower than Original. Yet large models show lower accuracy on this modified dataset: unlike counting individual mines, models must continuously track landmine cluster connectivity (e.g., judging cluster membership) for reasoning. This exceeds their simple pattern-matching capabilities, causing performance drops even for powerful models.
@@ 3.1 Experiment settings
On open-source models, we limit the reasoning budget to $32,768$ tokens before generating the final answer, regardless of their actual context window limitation. More specifically, we first input the prompt to the model to generate the reasoning part. If the model finishes reasoning within the budget, we then guide the model to generate the final answer that strictly follows the predefined JSON schema to eliminate presentation errors. A generation run is considered correct if and only if the model successfully finishes reasoning and produces a correct answer. We repeat $4$ runs on each sample with decoding temperature $T=0.6$. Closed-source models do not support hard reasoning budget limits, hence we simply limit their total output budget to $32,768$ tokens. Furthermore, we sample $5$ cases for each subtask due to expenditure constraints, while remaining repeating $4$ runs on each extracted sample. The prompt templates, including corresponding JSON schema for each game, are listed in Appendix C.2.
@@ 3.2 Main results
Figure 5 shows the comparison of the accuracy of each puzzle on both Original and the HardcoreLogic across all open-source models.***Due to the limited number of subtasks in some puzzles and the small sample size for testing such puzzles on closed-source models, all analysis of per game mainly focuses on open-source models. The overall performance of all puzzles and models shows a continuous downward trend. For open source models, Binario has the largest average performance degradation on the HardcoreLogic and Original. Skyscraper has the smallest decrease, followed by Navigation. These two puzzles are extremely difficult and extremely simple, which is why they have the smallest decrease.
@@ 4.1 Different long-tail transformation
In Section 2, we introduce four methods of long-tail transformation, including Search Space Expansion (IC1), Constraint Strengthening (IC2), Form Mutation (UE1), and Rule Mutation (UE2). Puzzles may also have two different long-tail transformation attributes at the same time. To quantify the impact of different long-tail transformations on puzzle difficulty, we fit a weighted multiple linear regression for each puzzle. The dependent variable represents the accuracy of the puzzle after undergoing four different long-tail transformations (IC1, IC2, UE1, UE2). We weight the number of samples in the data, and for mixed transformations, the predictive performance of the model is the sum of its coefficients. The intercept of the model represents the accuracy without any transformation, which is the average accuracy of each LRM on Original.
Figure 6 shows the coefficients of four long-tail transformations in the regression model, which is trained on data from open-source models and reflects their impact on puzzle difficulty. We can observe that IC1 has the greatest comprehensive impact on the models, as the increase in search space directly requires the improvement of the models’ memory and reasoning ability. UE1 requires the models to recognize some uncommon elements. It is worth noting that the parameter UE1 reaches its highest value for Sudoku puzzles, mainly due to the need to recognize irregular nine-grid patterns, indicating that the models struggles in this scenario. The parameters of the minesweeper puzzle in UE2 also show that the ”landmine cluster” rule has a significant impact on the models, which is consistent with our hypothesis in Section 2.
@@ 4.2 Error analysis
To probe the underlying causes of LRM failures on HardcoreLogic, we conduct a systematic error analysis. Based on the comparison between the puzzle, the correct answer, and the model’s complete responses, we identify six error categories: (1) Misunderstanding of the Logic Puzzle, (2) Misapplied Solution Framework, (3) Brute-Force with Excessive Complexity, (4) Factual Errors, (5) Faithfulness, and (6) Infinite Repetition. This enables us to move beyond aggregate accuracy in how different models fail. We randomly sample 50 erroneous cases from each of four representative models: gpt-oss-120b, the best-performing closed-source model on HardcoreLogic; Qwen3-235B, a representative of the Qwen series that we extensively evaluated; Kimi-K2-Instruct-Instruct, which experienced the largest performance drop from Original to HardcoreLogic; and Minimax-M1, the worst-performing model on HardcoreLogic. We employ GPT-5 (OpenAI, 2025b) as a secondary annotator to classify each case into one of the six categories. Detailed explanation of each category was shown in Appendix C.3.
@@
  • Factual errors are the most prevalent, suggesting that during extended reasoning, LRMs often fabricate facts to fill missing steps, compromising truthfulness and consistency. Mitigation may involve stronger penalties for factual deviations during fine-tuning or reinforcement learning, and mechanisms for intermediate reasoning verification.
@@ 4.3 Unsolvable games
To explore the model’s ability to handle contradictory puzzles, we constructed a batch of unsolvable puzzles based on each puzzle. Figure 8 shows the performance of each model in a puzzle-free scenario. It can be seen that the overall performance of closed-source models is better than that of open-source models. To investigate this phenomenon, we take several open-source models as examples to analyze the possible problems that models may encounter when encountering unsolvable puzzles.
@@
To investigate how LRMs handle unsolvable logic puzzles, we analyzed cases where LRMs correctly labeled puzzles as unsolvable to determine whether the judgment was a genuine understanding (Justified Unsolvability) or a heuristic claim due to failure to solve (Unjustified Unsolvability). We sampled responses from four models and classified each accordingly, revealing differences in their reasoning behavior, as shown in Figure 9. Stronger models (gpt-oss-120b and Qwen3-235B) typically provide justified explanations, while weaker models (Kimi-K2-Instruct and Minimax-M1) more often output unjustified “unsolvable” claims. This suggests that the ability to maintain deeper and more consistent reasoning chains is crucial for producing sufficient unsolvability explanations.

### v3

@@ Abstract
Large Reasoning Models (LRMs) have demonstrated impressive performance on complex tasks, including logical puzzle games that require deriving solutions satisfying all constraints. However, whether they can flexibly apply appropriate rules to varying conditions, particularly when faced with non-canonical game variants, remains an open question. Existing corpora focus on popular puzzles like 9x9 Sudoku, risking overfitting to canonical formats and memorization of solution patterns, which can mask deficiencies in understanding novel rules or adapting strategies to new variants. To address this, we introduce HardcoreLogic, a challenging benchmark of over 5,000 puzzles across 10 games, designed to test the robustness of LRMs on the “long-tail” of logical games. HardcoreLogic systematically transforms canonical puzzles through three dimensions: Increased Complexity (IC), Uncommon Elements (UE), and Unsolvable Puzzles (UP), reducing reliance on shortcut memorization. Evaluations on a diverse set of LRMs reveal significant performance drops, even for models achieving top scores on existing benchmarks, indicating heavy reliance on memorized stereotypes. While increased complexity is the dominant source of difficulty, models also struggle with subtle rule variations that do not necessarily increase puzzle difficulty. Our systematic error analysis on solvable and unsolvable puzzles further highlights gaps in genuine reasoning. Overall, HardcoreLogic exposes the limitations of current LRMs and establishes a benchmark for advancing high-level logical reasoning. Leaderboard: https://huggingface.co/spaces/JunsWan/HardcoreLogic Dataset: https://huggingface.co/datasets/xhWu-fd/HardcoreLogic Code: https://github.com/ljcleo/hardcore-logic
@@ 1 Introduction
Despite recent successes on benchmarks such as Enigmata (Chen et al., 2025a) and ZebraLogic (Lin et al., 2025a), whether LRMs are genuinely capable of true logical reasoning, i.e., flexibly apply appropriate rules to relevant conditions to derive correct conclusions, remains an important question. Take Sudoku as an example: while most real-world puzzles follow the canonical 9x9 format with nine 3x3 zones, variants with alternative constraints or irregular subgrids often prove challenging even for humans. Similarly, existing corpora exhibit a severe imbalance between canonical and non-canonical logic puzzles, making models prone to overfitting to canonical puzzles (Inger et al., 2025), leading to difficulties in solving non-canonical variants that fall into the long-tail of the distribution. This limitation manifests in two specific ways: (1) Models recognize only the canonical form of logical puzzles; when given a variant, they either struggle in understanding the new rules or ignore them, leading to faulty reasoning. (2) Models develop fixed solution strategies and reasoning patterns to solve canonical puzzles; even when they successfully understand the variant, they still apply a mismatched solution strategy, eventually producing errors or being confused.
@@
We further conduct a systematic error analysis to probe the underlying causes of model failures on HardcoreLogic. For solvable puzzles, we classify erroneous responses into six categories, and find that factual errors dominate across models, while more powerful models tend to exhibit brute-force errors, attempting exhaustive searches rather than strategic reasoning. Besides, models’ misunderstanding of problem constraints and misapplication of rigid rules lead to significant performance drops. For unsolvable puzzles, our analysis reveals that models performing well on solvable problems genuinely recognize unsolvability better. However, weaker models like Minimax-M1 may output "unsolvable" simply when they fail to find an answer, rather than through true recognition of logical unsatisfiability. When models fail to recognize unsolvability, we observe that stronger models mainly fail due to erroneous reasoning or inability to output answers within token budgets, while weaker models tend to force out solutions even without successfully deriving them. These highlight the need to improve models’ deep reasoning capabilities and robustness against degenerate behaviors. Overall, our contributions are threefold:
@@ 2.1 Preliminary: logical puzzle games
In HardcoreLogic, we focus on 10 types of logic games spanning 6 puzzle categories, including 8 challenging subtasks sourced from Enigmata (Chen et al., 2025a), the Zebralogic game from the ZebraLogic dataset (Lin et al., 2025a), and a classic Hanoi game synthesized by ourselves following its standard rules. All these three sources constitute the Original data used for comparison with HardcoreLogic. Specifically, HardcoreLogic covers the following 6 categories: (1) logic puzzle, (2) grid puzzle, (3) search puzzle, (4) pattern puzzle, (5) graph puzzleand (6) sequential puzzle. The 10 specific games are Zebralogic, Sudoku, Skyscraper, Kakurasu, Crypto, Navigation, Binario, Minesweeper, Hanoi and Hitori. See Appendix B.1 for a more detailed introduction.
@@ 2.2 Long-tail transformation
Standard logic puzzles are constrained in size, form diversity, and rule design, and thus fail to capture the irregularity and scale of real-world reasoning. To systematically construct more challenging evaluation data, we introduce a set of long-tail transformations that extend puzzles along three distinct dimensions: Increased Complexity, Uncommon Element, and Unsolvable Puzzle.
@@ 2.3 Complexity analysis
To systematically evaluate the hardness introduced by our long-tail transformations, we conceptualize complexity as a four-dimensional construct: Search Space Expansion (IC1), Constraint Strengthening (IC2), Form Variation (UE1), and Rule Variation (UE2). Each transformation is associated with dedicated quantitative metrics, and we compare all generated puzzles against the original benchmark. In the following, we present quantitative analyses of these four transformation types to demonstrate how each contributes to increased puzzle difficulty.
@@
This dimension captures the growth of candidate assignments induced by empty cells. Closed-form formulas are derived for each puzzle family (See Appendix B.4). For instance, in Binario, $N$ empty cells result in a search space of $\lvert S\rvert=2^{N}$. Figure 2 shows the average log-scale search space across five puzzle families, confirming that HardcoreLogic systematically enlarges the combinatorial space in all five games.
@@
  • For CSP-based puzzles (e.g., Zebralogic, Binario), we encode instances into Z3***Z3 refers to the Satisfiability Modulo Theories (SMT) solver developed by Microsoft Research. (de Moura & Bjørner, 2008) and collect: (i) Decisions: Explicit branching steps made by the solver; (ii) Conflicts: Backtracking events where partial assignments lead to contradictions. Larger counts indicate more complex search spaces and stronger constraint interactions, reflecting higher difficulty.
@@
Form variation (Ue1)
Form variation introduces novel symbols or forms that preserve puzzle validity but complicate comprehension. Since symbolic solvers cannot capture this representational difficulty, we measure it using perplexity, the inverse probability assigned by a pretrained LRM, which quantifies how surprising an instance appears. Higher perplexity values indicate that mutated forms impose greater representational complexity for LRMs. Figure 3 (light orange background) presents boxplots comparing the perplexity distributions of Original and HardcoreLogic instances; form variation consistently results in higher perplexity and thus greater representational difficulty.
Rule variation (Ue2)
Rule variation modifies or extends the logical rules governing puzzles, forcing solvers to adapt to new structural constraints.
@@
As shown in Figure 3 (light purple background), mutated-rule puzzles consistently yield higher solver statistics, indicating rule changes intensify reasoning complexity. A notable exception is the minesweeper dataset with "landmine clusters": numerical clues now represent adjacent landmine clusters, and more clues are added to ensure unique solutions—this reduces the search space, making required decisions lower than Original. Yet large models show lower accuracy on this modified dataset: unlike counting individual mines, models must continuously track landmine cluster connectivity (e.g., judging cluster membership) for reasoning. This exceeds their simple pattern-matching capabilities, causing performance drops even for powerful models.
@@ 3.1 Experiment settings
On open-source models, we limit the reasoning budget to $32,768$ tokens before generating the final answer, regardless of their actual context window limitation. More specifically, we first input the prompt to the model to generate the reasoning part. If the model finishes reasoning within the budget, we then guide the model to generate the final answer that strictly follows the predefined JSON schema to eliminate presentation errors. A generation run is considered correct if and only if the model successfully finishes reasoning and produces a correct answer. We repeat $4$ runs on each sample with decoding temperature $T=0.6$. Closed-source models do not support hard reasoning budget limits, hence we simply limit their total output budget to $32,768$ tokens. Furthermore, we sample $600$ cases across all games ($5$ per transformation type per game) due to expenditure constraints, while remaining repeating $4$ runs on each extracted sample. The prompt templates, including corresponding JSON schema for each game, are listed in Appendix C.2.
@@ 3.2 Main results
Figure 5 shows the comparison of the accuracy of each puzzle on both Original and the HardcoreLogic across all open-source models.Due to the limited number of subtasks in some puzzles and the small sample size for testing such puzzles on closed-source models, all analysis of per game mainly focuses on open-source models. The overall performance of all puzzles and models shows a continuous downward trend. For open source models, Binario has the largest average performance degradation on the HardcoreLogic and Original. Skyscraper has the smallest decrease, followed by Navigation. These two puzzles are extremely difficult and extremely simple, which is why they have the smallest decrease.
@@ 4.1 Different long-tail transformation
In Section 2, we introduce four methods of long-tail transformation, including Search Space Expansion (IC1), Constraint Strengthening (IC2), Form Variation (UE1), and Rule Variation (UE2). Puzzles may also have two different long-tail transformation attributes at the same time. To quantify the impact of different long-tail transformations on puzzle difficulty, we fit a weighted multiple linear regression for each puzzle. The dependent variable represents the accuracy of the puzzle after undergoing four different long-tail transformations (IC1, IC2, UE1, UE2). We weight the number of samples in the data, and for mixed transformations, the predictive performance of the model is the sum of its coefficients; the intercept represents the model-predicted baseline accuracy when no long-tail transformation is applied.
Figure 6 shows the coefficients of four long-tail transformations in the regression model, which is trained on data from open-source models and reflects their impact on puzzle difficulty. We can observe that IC1 has the greatest comprehensive impact on the models, as the increase in search space directly requires the improvement of the models’ memory and reasoning ability. UE1 requires the models to recognize some uncommon elements. It is worth noting that the parameter UE1 reaches its highest value for Sudoku puzzles, mainly due to the need to recognize irregular nine-grid patterns, indicating that the models struggles in this scenario. The parameters of the minesweeper puzzle in UE2 also show that the "landmine cluster" rule has a significant impact on the models, which is consistent with our hypothesis in Section 2.
@@ 4.2 Error analysis
To probe the underlying causes of LRM failures on HardcoreLogic, we conduct a systematic error analysis. Based on the comparison between the puzzle, the correct answer, and the model’s complete responses, we identify six error categories: (1) Misunderstanding of the Logic Puzzle, (2) Misapplied Solution Framework, (3) Brute-Force with Excessive Complexity, (4) Factual Errors, (5) Over Verfication, and (6) Infinite Repetition. This enables us to move beyond aggregate accuracy in how different models fail. We randomly sample 50 erroneous cases from each of four representative models: gpt-oss-120b, the best-performing closed-source model on HardcoreLogic; Qwen3-235B, a representative of the Qwen series that we extensively evaluated; Kimi-K2-Instruct, which experienced the largest performance drop from Original to HardcoreLogic; and Minimax-M1, the worst-performing model on HardcoreLogic. We employ GPT-5 (OpenAI, 2025b) as a secondary annotator to classify each case into one of the six categories. Detailed explanation of each category was shown in Appendix C.3.
We employ GPT-5 (OpenAI, 2025b), Gemini-2.5-Pro(Gemini Team, 2025), and Claude-Sonnet-4.5 (Anthropic, 2025) as secondary annotators to classify each case into one of the six categories. The final label is determined through a majority-vote scheme. In situations where the three models produce three distinct labels (i.e., no majority), we conduct manual verification. A detailed consistency analysis of this voting-and-adjudication scheme is provided in the Appendix  D.3.
@@
  • Factual errors are the most prevalent, suggesting that during extended reasoning, LRMs often fabricate facts to fill missing steps, compromising truthfulness and consistency. Mitigation may involve stronger penalties for factual deviations during fine-tuning or reinforcement learning, and mechanisms for intermediate reasoning verification. Meanwhile, during manual review, it was found that the Kimi-K2-Instruct’s significant skipping of steps during reasoning makes it more prone to introducing information that is not present in the problem or cannot be directly obtained during the reasoning process, thus making factual type errors more pronounced.
@@
To compare how model error patterns shift between Original and HardcoreLogic.we also uniformly sample 50 erroneous Original cases and classify them using the identical model-voting and human-verification pipeline. This provides a matched error-type distribution for Baseline errors, enabling a direct comparison against HardcoreLogic results. Figure 8 shows the percentage distribution of six error categories across the two benchmarks.
  • Rule Perturbation Raises Understanding-Related Failures HardcoreLogic introduces greater rule diversity, non-canonical puzzle structures, and more complex constraint dependencies. These perturbations substantially weaken models’ robustness in understanding and applying task rules. As a result, both Misunderstanding and Misapplied errors increase markedly across models.
  • Increased Complexity Reduces Plausible-but-Unfaithful Reasoning Over-Verification errors decline under HardcoreLogic, indicating that models are less able to generate coherent but incorrect explanations when faced with more complex logical dependencies. Instead of producing confident and polished but unfaithful rationales, models tend to break earlier in the reasoning process, yielding errors that stem from misunderstanding or rule misapplication.
Figure 8: Comparison of error-type distributions between Original and HardcoreLogic across four large language models.
@@ 4.3 Unsolvable games
To explore the model’s ability to handle contradictory puzzles, we constructed a batch of unsolvable puzzles based on each puzzle. Figure 9 shows the performance of each model in a puzzle-free scenario, where the overall performance of closed-source models is better than that of open-source models. To investigate this phenomenon, we take several open-source models as examples to analyze the possible problems that models may encounter when encountering unsolvable puzzles.
@@
To investigate how LRMs handle unsolvable logic puzzles, we analyzed cases where LRMs correctly labeled puzzles as unsolvable to determine whether the judgment was a genuine understanding (Justified Unsolvability) or a heuristic claim due to failure to solve (Unjustified Unsolvability). We sampled 50 responses from four models and classified each accordingly, revealing differences in their reasoning behavior, as shown in Figure 10. Stronger models (gpt-oss-120b and Qwen3-235B) typically provide justified explanations, while weaker models (Minimax-M1) more often output unjustified “unsolvable” claims. This suggests that the ability to maintain deeper and more consistent reasoning chains is crucial for producing sufficient unsolvability explanations.